OpenStudy (zuha):

Use the stem-and-leaf plot below to answer questions 1–3. 1. What is the range of data? 40 42 54 96 2. What is the median of the data? 64.5 72.5 74.5 75

OpenStudy (zuha):

here's the stem and leaf plot

OpenStudy (zuha):

3. What is the mode of the data? 5 70 75 88 i believe that the mode is 88

OpenStudy (zuha):

am I correct?

OpenStudy (zuha):

Use the stem-and-leaf plot below to answer questions 4–5. 4. How many students ran 6 or more miles? 7 5 4 3 5. What was the shortest distance, in miles, that a student ran? 4 4.3 43 43.5

OpenStudy (domebotnos):

Okay, you need to post them in separate questions, not all at once.

OpenStudy (domebotnos):

What are your answers for all of them?

OpenStudy (zuha):

here is the plot of Q 4-5

OpenStudy (zuha):

well I thought any one can help in any of the questions

OpenStudy (domebotnos):

What are your answers?

OpenStudy (zuha):

I'm still working on them

OpenStudy (domebotnos):

When you finish them give me your answers.

OpenStudy (zuha):

great ! can you help me solve them!

OpenStudy (domebotnos):

Do you think you know any of the answers?

OpenStudy (domebotnos):

To find the first answer, range is the highest number subtracted by the lowest number. 96-54 = ?

OpenStudy (zuha):

yes I just calculated it and it is 42

OpenStudy (domebotnos):

B.

OpenStudy (zuha):

let me check the 2nd one

OpenStudy (domebotnos):

For number 2. the middle number is the median. 74 and 75 are the middle numbers so solve. 74 + 75/2 =?

OpenStudy (domebotnos):

r u here?

OpenStudy (zuha):

yes, actually i was trying too :)

OpenStudy (zuha):

is it 111.5

OpenStudy (zuha):

no

OpenStudy (zuha):

let me check again

OpenStudy (zuha):

I'm not getting the correct answer !

OpenStudy (domebotnos):

74 + 75 = 149

OpenStudy (domebotnos):

149/2 = 74.5 :)

OpenStudy (zuha):

oh got it , 74.5

OpenStudy (zuha):

I was not calculating correctly

OpenStudy (zuha):

the next one must be 88?

OpenStudy (domebotnos):

3. Mode is the number that occurs the most often. The answer is NOT 88.

OpenStudy (zuha):

oh

OpenStudy (zuha):

75

OpenStudy (domebotnos):

Yes! :)

OpenStudy (zuha):

ok great

OpenStudy (domebotnos):

4. Count how many kids ran over 6 miles.

OpenStudy (zuha):

3

OpenStudy (zuha):

am I correct

OpenStudy (domebotnos):

No. Try again.

OpenStudy (zuha):

so should I go beyond 6up to 7 and up?

OpenStudy (domebotnos):

Count all the number above 6.

OpenStudy (domebotnos):

*numbers

OpenStudy (zuha):

ok

OpenStudy (zuha):

still confused !

OpenStudy (zuha):

is it 7

OpenStudy (domebotnos):

6.2, 6.2, 6.4, 7.1, 7.3, 8.0, 8.4 that is 7 numbers. Correct.

OpenStudy (zuha):

yes thats what i did

OpenStudy (domebotnos):

Good! :) Do u know what 5 is?

OpenStudy (zuha):

last Q is 43?

OpenStudy (domebotnos):

No, look at the key. (bottom right)

OpenStudy (zuha):

oh 4.3

OpenStudy (domebotnos):

Yes! All done! :D

OpenStudy (zuha):

thanks !! it was great studying wih ya:)

OpenStudy (domebotnos):

No problem! :D